Amedical researcher wishes to determine the percentage of females who take vitamins he wishes to be 99% confident that the estimate is within 2 percentage points of the true proportion a recent study of 180 females showed that 25% took vitamins. how large should the sample size be? if no estimate of the sample proportion is available how large should the sample be?
I do not know how to find what the sample size should be? so I am completely lost on this problem-I don't know where to start help!

Answer by stanbon(57323)

A medical researcher wishes to determine the percentage of females who take vitamins he wishes to be 99% confident that the estimate is within 2 percentage points of the true proportion a recent study of 180 females showed that 25% took vitamins. how large should the sample size be? if no estimate of the sample proportion is available how large should the sample be?
-------
Since E = z*sqrt(pq/n),
n = [z/E]^2*pq
---
Your 1st problem:
n = [2.5758/0.02]^2*0.25*0.75 = 3110
----
Your 2nd problem:
Use p = 1/2 as the best unbiased extimate.
n = [2.5758/0.02]^2 *0.5*0.5 = 4147 when rounded up
====================================================
Cheers,
Stan H.
=======================